# On the Birkhoff Spectrum for Hyperbolic Dynamics

### v1

@@ Abstract
In this paper, we study the structure of Birkhoff spectra for hyperbolic dynamical systems. For a Hölder observable $f$ on a basic set $\Lambda$, we prove that if the Birkhoff sums take both positive and negative values, then the Birkhoff spectrum $\mathcal{B}(f,\varphi,\Lambda)$ is dense in $\mathbb{R}$. This extends the density result of Gan, Shi, and Xia [6] from transitive Anosov diffeomorphisms in infranilmanifolds to general basic sets, and yields new density theorems for Axiom A systems, including Anosov diffeos.
Conversely, when the spectrum is not dense, we characterize concentrated observables, whose spectrum is confined to one side of zero. For these, we establish two rigidity results: (i) boundedness of the spectrum is equivalent to the function being cohomologous to a zero, which constitutes an extension of the Livšic theorem; (ii) if the spectrum exhibits an arithmetic structure, the function is cohomologous to a constant.
Finally, we extend the results to continuous time. For Anosov flows- including geodesic flows on Anosov manifolds-we obtain analogous density results for Birkhoff integrals along closed orbits. In particular, we generalize a theorem of Dairbekov–Sharafutdinov ([3]) by showing that a bounded or “arithmetically sparse” spectrum forces a smooth function to vanish or be constant, respectively.
@@ 1 Introduction
Our work extends and unifies several results in this direction, establishing density and rigidity properties for Birkhoff spectra on basic sets of diffeomorphisms and flows.
@@
We first extend a density theorem of [6] from Anosov diffeomorphisms in infranilmanifolds to general basic sets. The proof follows the same general strategy, but a key lemma ([6, Lemma 3.5]) is proved here using only the density of periodic measures in the space of invariant measures on a basic set, avoiding the geometric structure of infranilmanifolds used in the original reference.
Theorem 1.1.
Let $\Lambda\subset U$ be a basic set for $\varphi\colon U\to M$ and let $f\colon U\to\mathbb{R}$ be a Hölder continuous function. If $\mathcal{B}(f,\varphi,\Lambda)\cap\mathbb{R}^{+}\neq\emptyset$ and $\mathcal{B}(f,\varphi,\Lambda)\cap\mathbb{R}^{-}\neq\emptyset$, then $\mathcal{B}(f,\varphi,\Lambda)$ is dense in $\mathbb{R}$.
A continuous observable $f$ is called dispersed when
If instead $\mathcal{B}(f,\varphi,\Lambda)\subset\mathbb{R}_{\geq 0}$ or $\mathcal{B}(f,\varphi,\Lambda)\subset\mathbb{R}_{\leq 0}$, we say that $f$ is concentrated.
We immediately obtain consequences for Axiom A diffeomorphisms and transitive Anosov systems, which is a generalization of [6].
Corollary 1.2.
Let $\varphi$ be a $C^{1}$ Axiom A diffeomorphism and $f$ a dispersed Hölder continuous function. Then $\mathcal{B}(f,\varphi)$ is dense in $\mathbb{R}$. In particular, the result holds for Anosov diffeomorphisms, including transitive ones.
In the following theorems, we study Birkhoff sums for concentrated Hölder continuous functions on basic sets (when in general we do not have density). The most celebrated result in this direction is Livšic’s Theorem, which states (for basic sets) that if $\mathcal{B}(f,\varphi,\Lambda)=0$ for a Hölder continuous function $f$, then $f=u\circ\varphi-u$ for some Hölder continuous function $u$. In this case, we say that $f$ is cohomologous to zero.
Our next result we give a characterization to get some kind of density for concentrated functions. More specifically:
Theorem 1.3.
Let $\Lambda\subset U$ be a basic set and $f\colon U\to\mathbb{R}$ a concentrated Hölder continuous function. Then, $f$ is not cohomologous to zero and $\inf\mathcal{B}(f,\varphi,\Lambda)=0$ (resp. $\sup\mathcal{B}(f,\varphi,\Lambda)=0$) if and only if $\mathcal{B}(f,\varphi,\Lambda)$ is dense in $0,\infty)$ (resp. $(-\infty,0]$).
This result can be interpreted as a positivity induced denseness phenomenon: positivity (or negativity) of the observable, together with the infimum being zero, forces the range to be as large as possible-that is, dense in $0,\infty)$.
In other words, if a Hölder observable is nonnegative on a basic set and the infimum of its Birkhoff sums is zero, then the set of possible sums cannot be “sparse”; it must densely fill $0,\infty)$. The techinques developed in the proof of Theorem 1.1 and Theorem 1.3 can be used also study the set average Birkhoff Spectrum associate to a Hölder observable (see Section 3.2 for more details).
We characterize the boundedness of $\mathcal{B}(f,\varphi,\Lambda)$ for concentrated Hölder continuous functions, which extends Livšic’s theorem.
@@
Theorem 1.4 establishes a sharp rigidity principle for Hölder observables on basic sets: boundedness of the Birkhoff range $\mathcal{B}(f,\varphi,\Lambda)$ forces the observable to be cohomologous to a constant. This equivalence turns a mild, qualitative condition—boundedness—into a powerful geometric constraint, showing that the only way the set of Birkhoff integrals can avoid being unbounded is if all invariant measures yield the same average. In particular, it extends Livšic’s theorem from the zero-average case to arbitrary constants and reveals a fundamental dichotomy: either $\mathcal{B}(f,\varphi,\Lambda)$ is a singleton, or it is unbounded. This result lies at the heart of the geometric theory of hyperbolic dynamics, with direct implications for thermodynamic formalism, and the multifractal analysis of Birkhoff spectra.
@@
The methods developed for proving Theorem 1.4 yield deeper rigidity phenomena, revealing how arithmetic constraints on the Birkhoff range force the observable to be trivial. While Theorem 1.4 shows that boundedness of $\mathcal{B}(f,\varphi,\Lambda)$ implies the function is cohomologous to a constant, the following result demonstrates that a different kind of restriction-namely, the set being contained in a finite union of dilated copies of $\mathbb{N}$-already forces similar conclusion. This can be interpreted as a quantization rigidity property: the only way for the set of Birkhoff averages to be “arithmetically sparse” is if all invariant measures give the same average.
@@
Assume that $\mathcal{B}(f,\varphi,\Lambda)\subset a_{1}\mathbb{N}\cup\dots\cup a_{l}\mathbb{N}$ for some $a_{1},\dots,a_{l}\in\mathbb{R}$. Then $f$ is cohomologous to a constant $a_{i_{0}}$ for some $i_{0}$.
@@
Let $\phi^{t}\colon U\to N$ be a flow, $\Lambda\subset U$ a basic set for $\phi^{t}$, and $f\colon U\to\mathbb{R}$ a dispersed Hölder continuous function. Then $\mathcal{B}(f,\phi^{t})$ is dense in $\mathbb{R}$. In particular, the result holds for Anosov flows, including transitive ones.
When $f$ is not dispersed, results analogous to Theorems 1.3, 1.4, and 1.5, are obtained in the setting of Anosov flows.
@@
In contrast with Theorems 1.4 and 1.3 for the flow case, the geodesic flow exhibits a fundamentally different behavior. As shown in Section 4.3 (see Remark 4.6), for the specific choice $f=\text{Ric}$ (the Ricci curvature), the associated $\mathcal{B}(\mathrm{Ric},\phi^{t})$ is unbounded, and 0 is not an accumulation point of its spectrum.
@@
Let $(M,g)$ be an Anosov manifold with geodesic flow $\phi^{t}_{M}$. For every dispersed Hölder function $f$,
@@
    If $\mathcal{B}(f,\varphi^{t})\subset a_{1}\mathbb{N}\cup\dots\cup a_{l}\mathbb{N}$ for some $a_{1},\dots,a_{l}\in\mathbb{R}$. Then $f$ is constant $a_{i_{0}}$ for some $i_{0}$.
@@
The paper is organized as follows. Section 2 recalls necessary background on hyperbolic dynamics, periodic measures, and the Livšic Theorem. Section 3 contains the proofs of our main theorems for diffeomorphisms, including an application to the average Birkhoff spectrum (Theorem 3.7). Section 4 extends these results to flows. After reducing the problem via a Poincar’e map in Section 4.1, we prove Theorems 1.6 and 1.8. We conclude with applications to geodesic flows in Section 4.3.
@@ 2.1 Hyperbolic sets and basic sets
A continuous real-valued function $f$ on $\Omega$ is called dispersed if there exists an index $i_{0}\in{1,\dots,k}$ such that: $\mathcal{B}(f,\varphi,\Omega_{{i_{0}}})\cap\mathbb{R}^{\pm}\neq\emptyset$. Otherwise, the function $f$ is said to be concentrated. The density in the theorems of the last section will be achieved only in the dispersed function of Hölder.
@@ 2.4 Livšic Theorem and cohomology
This rigidity plays a central role in our analysis of concentrated functions (Theorems 1.4 and 1.5).
@@ 3 Proof of Main Results for Diffeomorphisms
3.1 Proof of Theorem 1.1
For our purpose, we recall the concept of asymptotically rational independence of a sequence of real numbers [6] and [7].
@@
    If condition (1) fails for all pairs of periodic points, we prove that there exists a sequence $\{p_{n}\}$ converging to $p$ such that $S(f,\varphi,p_{n})$ is asymptotically rationally independent of $S(f,\varphi,q)$. This step represents a crucial difference from [6], since we do not rely on the structure of infranilmanifolds to ensure this. Instead, we use the density of periodic measures in the space of invariant measures for a basic set (Sigmund’s theorem).
@@
To guarantee step (2), we invoke the following result of Sigmund on the density of periodic invariant probability measures:
Theorem 3.3 ([17]).
If $\Lambda$ is a basic set for a diffeomorphism $\varphi$, then the set
We will use this result to prove the following lemma (compare with [6, Lemma 3.5]).
@@
Let $\Lambda$ be a basic set for a diffeomorphism $\varphi$, and let $f\colon\Lambda\to\mathbb{R}$ be a continuous function and $S(f,\varphi,p)<0<S(f,\varphi,q)$ for some $p,q\in\textrm{Per}(\varphi)$. If for every $z\in\textrm{Per}(\varphi)$, the sums $S(f,\varphi,z)$ and $S(f,\varphi,q)$ are rationally dependent, then there exists a sequence $p_{n}\in\textrm{Per}(\varphi)$ such that $S(f,\varphi,p_{n})$ is asymptotically rationally independent of $S(f,\varphi,q)$.
@@
It is not difficult to prove that the pair $S(f,\varphi,z)$ and $S(f,\varphi,q)$ are rationally dependent if and only if the normalized sums $\frac{S(f,\varphi,z)}{\tau_{z}}$ and $\frac{S(f,\varphi,q)}{\tau_{q}}$ are rationally dependent. Therefore, it is sufficient to prove the lemma for the normalized sums.
@@
$k_{n}\cdot\frac{S(f,\varphi,p_{n})}{\tau_{n}}-l_{n}\cdot\frac{S(f,\varphi,q)}{\tau_{q}}=0.$
Define $a_{n}:=\frac{S(f,\varphi,p_{n})}{\tau_{n}}$ and $b:=\frac{S(f,\varphi,q)}{\tau_{q}}\neq 0$.
@@
$\left\{\frac{S(f,\varphi,p_{n})}{\tau_{p_{n}}}:n\geq 1\right\}\subset c\cdot\mathbb{Z}.$
From Theorem 3.3, for any $\varphi$-invariant probability measure $\mu$, there exists a subsequence $(p_{n_{k}})$ such that
$\frac{S(f,\varphi,p_{n_{k}})}{\tau_{p_{n_{k}}}}\to\int f\,d\mu.$
Consequently, $\int f\,d\mu\in c\cdot\mathbb{Z}$ for all $\mu\in\mathcal{M}(\Lambda,\varphi)$. Since $S(f,\varphi,p)<0<S(f,\varphi,q)$, then we can find two different invariant probability measures, $\mu_{1}$ and $\mu_{2}$, such that $\int f\,d\mu_{1}\neq\int f\,d\mu_{2}$.
Define a one-parameter family of invariant measure $\mu_{t}=t\mu_{1}+(1-t)\mu_{2}$ for $t\in[0,1]$. Then,
$\int f\,d\mu_{t}=t\int f\,d\mu_{1}+(1-t)\int f\,d\mu_{2}\in c\cdot\mathbb{Z}\quad\text{for all }t\in[0,1].$ (3)
However, this leads to a contradiction. Since $\int f\,d\mu_{1}$ and $\int f\,d\mu_{2}$ are fixed elements of $c\cdot\mathbb{Z}$, the linear combination in (3) can only yield another element of $c\cdot\mathbb{Z}$ for at most finitely many parameters $t\in[0,1]$, not for all of them. This contradiction completes the proof of the lemma. ∎
Remark 3.5.
It is worth noting that Lemma 3.4 is more general than [6, Lemma 3.5], as it does not require the manifold to be an infranilmanifold. This generalization allows us to extend [6, Theorem 1.1] to more general cases, such as Anosov diffeomorphisms on arbitrary manifolds without any transitivity conditions (see Corollary 1.2). So, this remark will be very useful in understanding some equivalences between Birkhoff spectrum and average Birkhoff spectrum for basic sets.
Remark 3.6.
In the end of the proof Lemma 3.4, we used Theorem 3.3 to obtain a contradiction. But it is great to observe that it is enough assuming that the set
$\Big\{\frac{S(f,\varphi,p)}{\tau_{p}}:p\in\textrm{Per}(\varphi)\Big\}\not\subset c\mathbb{Z},\,\,\text{for any}\,\,c\in\mathbb{R}.$
@@
From our hypotheses, assume there exist $p,q\in\textrm{Per}(\varphi)$ such that
@@
Case 2: For every $z\in\textrm{Per}(\varphi)$, the values $S(f,\varphi,z)$ and $S(f,\varphi,q)$ are rationally dependent. In this case, applying Lemma 3.4 together with [6, Lemma 3.3 and Remark 3.4] completes the proof of the theorem. ∎
Proof of Corollary 1.2.
It is just using the Smale’s Spectral Decomposition Theorem and Theorem 1.1.

### v2

@@ Abstract
In this paper, we study the structure of Birkhoff spectra for hyperbolic dynamical systems. Given a Hölder observable $f$ on a basic set $\Lambda$, we obtain the following results: First, we characterize when the Birkhoff spectrum of $f$ is dense in the positive (or negative) real line. Second, we prove that a bounded Birkhoff spectrum forces $f$ to be cohomologous to zero, which constitutes an extension of Livšic’s theorem. Moreover, we show that if the spectrum exhibits an “arithmetically sparse” structure, then $f$ is cohomologous to a constant.
We then extend these results to continuous time. For Anosov flows—including geodesic flows on Anosov manifolds—we establish analogous density results for Birkhoff integrals over closed orbits. In particular, we generalize a theorem of Dairbekov–Sharafutdinov [3] by proving that a bounded (resp. arithmetically sparse) spectrum forces a smooth function to vanish (resp. be constant).
@@ 1 Introduction
This work, inspired by [6], explores the density of Birkhoff spectra for some Hölder continuous observables, establishing new rigidity results for these spectra on basic sets of diffeomorphisms and flows.
@@
A celebrated result, Livšic’s Theorem [9], characterizes when a Hölder continuous function $f$ has a trivial Birkhoff spectrum on a basic set, i.e., when $\mathcal{B}(f,\varphi,\Lambda)=0$. It establishes that this occurs if and only if $f$ is cohomologous to a constant—specifically, if there exists a Hölder continuous function $u$ such that $f=u\circ\varphi-u$. In this case, we say that $f$ is cohomologous to zero.
Our first result establishes a sharp rigidity principle for Hölder observables on basic sets: the boundedness of the Birkhoff spectra $\mathcal{B}(f,\varphi,\Lambda)$ forces the observable to be cohomologous to a constant. More specifically,
@@
This equivalence turns a mild, qualitative condition—boundedness—into a powerful geometric constraint, showing that the only way the set of Birkhoff integrals can avoid being unbounded is if all invariant measures yield the same average. In particular, it extends Livšic’s theorem from trivial spectrum case to arbitrary boundedness condition and reveals a fundamental dichotomy: either $\mathcal{B}(f,\varphi,\Lambda)$ is a singleton, or it is unbounded. This result lies at the heart of the geometric theory of hyperbolic dynamics, with direct implications for thermodynamic formalism, and the multifractal analysis of Birkhoff spectra.
@@
To state our second result, we introduce a classification of observables based on their spectral support. A continuous observable $f$ is termed dispersed if its Birkhoff spectrum intersects both the positive and negative reals, i.e.,
Conversely, if $\mathcal{B}(f,\varphi,\Lambda)\subset\mathbb{R}_{\geq 0}$ or $\mathcal{B}(f,\varphi,\Lambda)\subset\mathbb{R}_{\leq 0}$, we call $f$ concentrated.
Moreover, a continuous observable $f$ is called arithmetic if there is $c\in\mathbb{R}$ with
$\mathcal{B}(f,\varphi,\Lambda)\subset c\mathbb{Z}.$
In [6], it was shown that for transitive Anosov diffeomorphisms and any dispersed observable, the Birkhoff spectrum is dense in $\mathbb{R}$. In the same paper, the density result was obtained for basic sets (instead of Anosov systems) for dispersed and non-arithmetic observable (see Theorem 3.1 in Section 3.1 for details). However, in Section 3.2, we provide an example of a basic set and a dispersed non-arithmetic observable for which density of the Birkhoff spectra fails. This suggests that a sufficient condition for density likely involves additional topological restrictions on the basic set. The techniques developed in the proofs of Theorem 3.1 and Theorem 1.1 can also be used to study the average Birkhoff spectrum associated with a Hölder observable (see Section 3.3 for details).
Concentrated observables preclude spectral density in $\mathbb{R}$, our subsequent result provides sufficient conditions to achieve density in the maximal admissible subset.
@@
Let $\Lambda\subset U$ be a basic set for $\varphi\colon U\to M$ and let $f\colon U\to\mathbb{R}$ be a Hölder continuous concentrated function. Then, if $f$ is not cohomologous to zero and $\inf\mathcal{B}(f,\varphi,\Lambda)=0$ (resp. $\sup\mathcal{B}(f,\varphi,\Lambda)=0$) if and only if $\mathcal{B}(f,\varphi,\Lambda)$ is dense in $0,\infty)$ (resp. $(-\infty,0]$).
This result can be interpreted as a concentrated induced denseness phenomenon: positivity (or negativity) of the observable, together with the infimum (or maximum)being zero, forces the range to be as large as possible-that is, dense in $0,\infty)$ (or $(-\infty,0]$).
While Theorem 1.1 shows that boundedness of $\mathcal{B}(f,\varphi,\Lambda)$ implies that $f$ is cohomologous to zero, the following result demonstrates that a significantly weaker restriction-namely, the set being contained in a finite union of dilated copies of $\mathbb{Z}$-forces the similar conclusion. This can be interpreted as a quantization rigidity property: the only way for the Birkhoff spectra to be “arithmetically sparse” is if all Birkhoff averages coincide.
Theorem 1.3.
Let $\Lambda\subset U$ be a basic set for $\varphi\colon U\to M$ and let $f$ be a Hölder continuous non-arithmetic observable. If
$\mathcal{B}(f,\varphi,\Lambda)\subset a_{1}\mathbb{Z}\cup\dots\cup a_{l}\mathbb{Z}\quad\text{for some }a_{1},\dots,a_{l}\in\mathbb{R}^{+},$
then $f$ is cohomologous to a constant $a_{i_{0}}$ for some $i_{0}$.
The non-arithmeticity condition on $f$ is necessary, as shown in Section 3.2.
@@
Let $\phi^{t}\colon U\to N$ be a flow, $\Lambda\subset U$ a basic set for $\phi^{t}$, and $f\colon U\to\mathbb{R}$ a dispersed and non-arithmetic Hölder continuous function. Then $\mathcal{B}(f,\phi^{t})$ is dense in $\mathbb{R}$. In particular, for Anosov flows (including transitive ones) the conclusion holds without requiring the non-arithmeticity condition.
When $f$ is not dispersed, results analogous to Theorems 1.1, 1.2, and 1.3, are obtained in the setting of Anosov flows.
@@
In contrast with Theorems 1.1 and 1.2 for the flow case, the geodesic flow exhibits a fundamentally different behavior. As shown in Section 4.3 (see Remark 4.6), for the specific choice $f=\text{Ric}$ (the Ricci curvature), the associated $\mathcal{B}(\mathrm{Ric},\phi^{t})$ is unbounded, and 0 is not an accumulation point of its spectrum.
@@
Let $(M,g)$ be an Anosov manifold with geodesic flow $\phi^{t}_{M}$. For every dispersed and non-arithmetic Hölder function $f$,
@@
    If $f$ is non-arithmetic and $\mathcal{B}(f,\varphi^{t})\subset a_{1}\mathbb{Z}\cup\dots\cup a_{l}\mathbb{Z}$ for some $a_{1},\dots,a_{l}\in\mathbb{R}$. Then $f$ is constant $a_{i_{0}}$ for some $i_{0}$.
@@
The paper is organized as follows. Section 2 recalls necessary background on hyperbolic dynamics, periodic measures, and the Livšic Theorem. Section 3 contains the proofs of our main theorems for diffeomorphisms, including an application to the average Birkhoff spectrum (Theorem 3.6). Section 4 extends these results to flows. After reducing the problem via a Poincar’e map in Section 4.1, we prove Theorems 1.4 and 1.6. We conclude with applications to geodesic flows in Section 4.3.
@@ 2.1 Hyperbolic sets and basic sets
A continuous real-valued function $f$ on $\Omega$ is called dispersed if there exists an index $i_{0}\in{1,\dots,k}$ such that: $\mathcal{B}(f,\varphi,\Omega_{{i_{0}}})\cap\mathbb{R}^{\pm}\neq\emptyset$. Otherwise, the function $f$ is said to be concentrated.
Moreover, a continuous observable $f$ is called arithmetic if for all index $i\in{1,\dots,k}$ there is $c_{i}\in\mathbb{R}$ with
$\mathcal{B}(f,\varphi,\Omega_{{i}})\subset c_{i}\mathbb{Z}.$
The density in the theorems of the last section will be achieved only in the dispersed and non-arithmetic function of Hölder.
@@ 2.4 Livšic Theorem and cohomology
This rigidity plays a central role in our analysis of concentrated functions (Theorems 1.1 and 1.3).
@@ 3 Proof of Main Results for Diffeomorphisms
3.1 Density of the Birkhoff Spectra
In this section, we establish conditions that guarantee density of the spectra for basic sets. This result was obtained in [6]; however, our exposition aims to clarify the origin of the ”non-arithmeticity” condition. More precisely,
Theorem 3.1 ([6]).
Let $\Lambda\subset U$ be a basic set for $\varphi\colon U\to M$ and let $f\colon U\to\mathbb{R}$ be a Hölder continuous function. If $f$ is dispersed and non-arithmetic, then $\mathcal{B}(f,\varphi,\Lambda)$ is dense in $\mathbb{R}$. In particular, for Anosov diffeomorphisms (including transitive ones) the conclusion holds without requiring the non-arithmeticity condition.
The proof of this result uses some concepts we present below. The first is the notion of asymptotically rational independence for sequences of real numbers, as introduced in [6] and [7].
@@ 3.1 Density of the Birkhoff Spectra
    If condition (1) fails for all pairs of periodic points, we show that there exists a sequence $\{p_{n}\}$ converging to $p$ such that $S(f,\varphi,p_{n})$ is asymptotically rationally independent of $S(f,\varphi,q)$. If no such sequence exists, then Lemma 3.3 implies that $f$ is arithmetic, yielding a contradiction.
@@
[6, Lemma 3.5] Let $\Lambda$ be a basic set for a diffeomorphism $\varphi$, and let $f\colon\Lambda\to\mathbb{R}$ be a continuous function which is non-arithmetic. If for every $z\in\textrm{Per}(\varphi)$, the sums $S(f,\varphi,z)$ and $S(f,\varphi,q)$ are rationally dependent, then there exists a sequence $p_{n}\in\textrm{Per}(\varphi)$ such that $S(f,\varphi,p_{n})$ is asymptotically rationally independent of $S(f,\varphi,q)$.
@@
$k_{n}\cdot S(f,\varphi,p_{n})-l_{n}\cdot S(f,\varphi,q)=0.$
Define $a_{n}:=S(f,\varphi,p_{n})$ and $b:=S(f,\varphi,q)\neq 0$.
@@
$\left\{S(f,\varphi,p_{n}):n\geq 1\right\}\subset c\cdot\mathbb{Z}.$
This implies that $f$ is arithmetic, which is a contradiction. ∎
@@
From our hypotheses, since $f$ is dispersed, assume there exist $p,q\in\textrm{Per}(\varphi)$ such that
@@
Case 2: For every $z\in\textrm{Per}(\varphi)$, the values $S(f,\varphi,z)$ and $S(f,\varphi,q)$ are rationally dependent. In this case, applying Lemma 3.1 together with [6, Lemma 3.3 and Remark 3.4] completes the proof of the theorem. Finally, in the case of a transitive Anosov diffeomorphism, if $f$ is arithmetic, then $f$ is cohomologous to a constant and hence concentrated. This yields a contradiction. Therefore, for Anosov systems, the dispersed condition alone suffices to guarantee density (see [6, Proof of Lemma 3.5]). ∎
We immediately obtain consequences for Axiom A diffeomorphisms.
Corollary 3.5.
Let $\varphi$ be a $C^{1}$ Axiom A diffeomorphism and $f$ a dispersed and non-arithmetic Hölder continuous function. Then $\mathcal{B}(f,\varphi)$ is dense in $\mathbb{R}$. In particular, the result holds for Anosov diffeomorphisms. In particular, for Anosov diffeomorphisms (including transitive ones) the conclusion holds without requiring the non-arithmeticity condition.
Proof.
It is just using the Smale’s Spectral Decomposition Theorem and Theorem 3.1. ∎
3.2 Example of dispersed and arithmetic Lipchitz Function on basic set
The following example demonstrates that the non-arithmeticity condition is necessary for the density conclusion in Theorem 3.1. Thus, the only way to obtain density solely from the dispersed condition is when $\Lambda$ has additional topological properties Let $\Lambda=\{0,1\}^{\mathbb{Z}}$ with shift $\sigma$, and equip $\Lambda$ with the metric
$d(x,y)=\sum_{k\in\mathbb{Z}}2^{-|k|}|x_{k}-y_{k}|.$
Consider the function $f\colon\Lambda\to\mathbb{R}$ define by $f(x)=x_{0}+x_{1}-1$, where $x=(x_{n})_{n\in\mathbb{Z}}$.
The function $f$ has the following properties
  1. 1.
    $f$ is Lipschitz continuity (in particular Hölder).
    For any $x,y\in\Lambda$:
    $|\varphi(x)-\varphi(y)|$ $=$ $|(x_{0}+x_{1}-1)-(y_{0}+y_{1}-1)|=|(x_{0}-y_{0})+(x_{1}-y_{1})|$
    $\leq$ $d(x,y)+2d(x,y)=3d(x,y).$
  2. 2.
    Not cohomologous to any constant and dispersed.
    Suppose, for contradiction, that $\varphi$ is cohomologous to a constant $c\in\mathbb{R}$. That is, there exists a continuous function $\eta:\Lambda\to\mathbb{R}$ such that
    $f(x)=c+\eta(\sigma(x))-\eta(x)\quad\text{for all }x\in\Lambda.$
    Consider the two fixed points:
    $e_{0}=(\dots,0,0,0,\dots),\quad e_{1}=(\dots,1,1,1,\dots).$
    Observe that $f(e_{0})=-1$, and $f(e_{1})=1$. Therefore, $f$ is dispersed.
    Since $\sigma(e_{0})=e_{0}$ and $\sigma(e_{1})=e_{1}$, then the coboundary condition gives:
    $1=c+\eta(e_{0})-\eta(e_{0})=c\,\,\,\text{and}\,\,\,1=c+\eta(e_{1})-\eta(e_{1})=c$
    Thus, we obtain c=1=1c=-1=1, a contradiction. Therefore $\varphi$ cannot be cohomologous to any constant $c\in\mathbb{R}$.
  3. 3.
    $f$ is arithmetic.
    Let $x$ be a periodic point with minimal period $p\geq 1$ and repeating block
    $(a_{0},a_{1},\dots,a_{p-1}),\quad a_{k}\in\{0,1\}.$
    That is, $x_{k}=a_{k\bmod p}$ for all $k\in\mathbb{Z}$.
    The Birkhoff sum over one period is:
    $S(f,\sigma,x)$ $=$ $\sum_{k=0}^{p-1}f(\sigma^{k}x)=\sum_{k=0}^{p-1}\big(a_{k}+a_{k+1\bmod p}-1\big)$
    $=$ $2\sum_{k=0}^{p-1}a_{k}-p\in\mathbb{Z},$
    since $\sum_{k=0}^{p-1}a_{k+1\bmod p}=\sum_{k=0}^{p-1}a_{k}.$